### v1

# Drawback of Enforcing Equivariance and its Compensation via the Lens of Expressive Power

Abstract
Equivariant neural networks encode symmetry as an inductive bias and have achieved strong empirical performance in wide domains. However, their expressive power remains not well understood. Focusing on 2-layer ReLU networks, this paper investigates the impact of equivariance constraints on the expressivity of equivariant and layer-wise equivariant networks. By examining the boundary hyperplanes and the channel vectors of ReLU networks, we construct an example showing that equivariance constraints could strictly limit expressive power. However, we demonstrate that this drawback can be compensated via enlarging the model size. Furthermore, we show that despite a larger model size, the resulting architecture could still correspond to a hypothesis space with lower complexity, implying superior generalizability for equivariant networks.
1 Introduction
In scaling up the capacity of machine learning models, leveraging the intrinsic symmetry in data, if it exists, would be a much more efficient approach than simply scaling up model size. Good examples of this intrinsic symmetry would be point cloud processing qi2017pointnet; li2018pointcnn; fuchs2020se; chen2021equivariant, graph neural networks velivckovic2017graph; kanatsoulis2024graph, and chemistry (faber2016machine; eismann2021hierarchical). When learning these symmetries, vanilla neural network-based methods must be shown numerous (sometimes manually) transformed data examples, such as rotated or translated images, and enforced to make invariant/equivariant predictions, which is usually computationally expensive and requires vast amounts of data (bronstein2021geometric). Equivariant neural networks overcome these challenges by constraining their output to be invariant or equivariant to group representations (cohen2016group), through either new architectures or optimization methods. One useful approach is to enforce each layer of the network to be equivariant.
While the equivariance constraint in a neural network guarantees the symmetry properties, it remains unclear whether they compromise the expressive power. In this paper, we address this by determining whether equivariance (or layer-wise equivariance) increases the expected risk when approximating an equivariant target function. Specifically, we examine if the following strict inequality holds for some equivariant target function $s$ and some parameter spaces $\Theta$:
$\inf_{\theta\in\Theta}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}]$ $<\inf_{\theta\in\Theta\cap\text{GENs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}]$
$\text{and}~~~~\inf_{\theta\in\Theta\cap\text{GENs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}]$ $<\inf_{\theta\in\Theta\cap\text{LENs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}].$
where $\Theta$ is the set of network parameters (usually $L$ matrices) and $F_{\theta}$ is the neural network with parameter $\theta$. Notation GENs means the model parameters that make $F_{\theta}$ an equivariant network, while LENs is the parameter set of layer-wise equivariant networks.
Figure 1: An equivariant function that satisfies $s((a,b)^{\top})=s((b,a)^{\top})$ and $s((a,b)^{\top})=s((-a,-b)^{\top})$. The left subfigure is a 3D plot, while the right subfigure is a 2D Contour map.
Interpretation of Equivariant Constraints.
The equivariance constraints primarily govern the input-output mapping. To understand how this affects the model parameters, we focus on the required weight matrices. We demonstrate that for 2-layer ReLU networks, requiring equivariance forces the network to have specific structural symmetries. For GENs, this necessity manifests as the boundary hyperplanes that define the network’s linear regions must be symmetric. While the layer-wise equivariance constraint (LENs) requires sufficient model capacity to contain symmetric channel vectors. This framework, which connects the functional requirement of equivariance to its necessary structural consequences, provides a clear basis for comparing their expressive power.
Drawback of layer-wise equivariance in expressive power.
Following the insight of how equivariance restricts model parameters, we demonstrate a scenario where equivariance strictly hurts the expressive power. Specifically, we construct an invariant target function $s$ on $\mathbb{R}^{2}$ and establish two key findings: 1) for a 1-neuron 2-layer architecture, networks without equivariance constraints achieve a strictly lower expected error compared to GENs, and 2) for a 2-layer architecture, LENs require a strictly larger model size than GENs to fully approximate the target function.
Compensations via enlarging model size.
Inversely, we prove that an increase in model size can fix the reduced expressive power caused by (layer-wise) equivariance. When the target function is invariant, and the data distribution is symmetric, enlarging the model size of any equivariant network by $|G|$ times can compensate for the expected loss over a 2-layer neural network. Moreover, a doubled model size allows LENs to represent all invariant functions of GENs. We further construct a LEN with $|G|$ times enlarged model size whose hypothesis space has a lower dimension. This implies that the LEN has improved generalizability; together with the comparable expressive power, this result renders LENs a competitive solution.
To the best of our knowledge, this paper provides the first in-depth theoretical examination of the expressive power of layer-wise equivariant neural networks. We also propose that an appropriately larger model size can compensate for the loss in expressive power and potentially enhance model generalisability, underlining the advantages of incorporating layer-wise equivariance into model architectures.
2 Related Works
Designing equivariant neural networks. Remarkable advances have been seen in designing equivariant network architectures for the symmetries in data. Typical approaches for obtaining an equivariant neural network can be designing equivariant architectures cohen2016steerable; satorras2021n or introducing equivariance constraints into the optimization winter2022unsupervised; tang2022equivariance.
Equivariant neural architectures. A typical designing paradigm is designing equivariant neural architectures based on classic neural networks. For example, cohen2016group design group convolutional neural networks in the group spaces as the equivariant variants of the convolutional neural networks shin2016deep, which is extended to the homogeneous spaces by cohen2019general. Similar approaches are also seen in designing equivariant graph neural networks klicpera2020directional, equivariant transformers hutchinson2021lietransformer, equivariant diffusion hoogeboom2022equivariant, equivariant point network chen2021equivariant, equivariant harmonic networks worrall2017harmonic, etc.
Symmetry-aware optimization. In this stream, equivariance constraints on the the parameters of the networks are introduced in the optimization to ensure the equivariance of the learned model, usually called steerable neural networks cohen2016steerable. For examples, methods surge to compute the solution space satisfying the equivariance constraints for different groups and network architectures, including the symmetry group $S_{N}$ in steerable graph networks maron2018invariant and the rotation and Euclidean groups $\{C_{N},E_{2}\}$ in steerable convolutional neural networks weiler2018learning; weiler2019general. A major obstacle is the high computational costs to compute the equivariant basis. finzi2021practical address this issues with an algorithm with polynomial computational complexity on the sum of the number of discrete generators and the dimension of the group, which divides the problem into some independent subproblems and adopts Krylov method to compute nullspaces.
Theoretical analyses. With many works contributing in deep learning theory he2020recent, theoretical studies have shown that equivariant neural networks have significant advantages from the aspects of generalizability, convergence, and approximation. sannai2021improved prove improved generalization error bounds of equivariant models . lawrence2021implicit prove that linear group convolutional neural networks trained by gradient descent for binary classification converge to solutions with low-rank Fourier matrix coefficients based on the results via implicit bias. zaheer2017deep; maron2019universality; yarotsky2021universal prove the universal approximation ability of equivariant models by various approximation techniques, while ravanbakhsh2020universal utilize group averaging to obtain an equivariant approximator to prove the universal approximation ability. In addition, elesedy2021provably prove that invariant/equivariant models have a smaller expected loss when the target function is invariant/equivariant. azizian2021expressive study the expressive power of invariant/equivariant graph neural networks.
3 Problem Settings
In this section, we introduce our problem setting. We focus on two-layer ReLU networks with invariant output functions to show how equivalence and layer-wise equivalence impact the expressive power. Besides, we also assume the groups to be finite.
General networks (GNs)
GNs are neural networks without other constraints. In this paper, we focus on the architecture of a two-layer network, which forms a function $F:\mathbb{R}^{n}\to\mathbb{R}^{d}$. The network consists of a hidden layer with weight matrix $W^{(1)}\in\mathbb{R}^{n\times m}$ and an output layer with weight matrix $W^{(2)}\in\mathbb{R}^{m\times d}$ where $m$ is the size of the hidden layer and the number of neurons. With ReLU nonlinearity $\sigma$, the input $x\in\mathbb{R}^{n}$ is mapped to the output space by $F(x)=W^{(2)}\sigma(W^{(1)}x)$. For simplicity, we represent $\alpha_{i}^{\top}$ as the $i$-th column of $W^{(1)}$ and $\beta_{j}$ as the $j$-th row of $W^{(2)}$ and $W^{(1)}$. Then the output of a GN can be formulated as follows:
$F(x)=W^{(2)}\sigma(W^{(1)}x)=\sum_{i=1}^{m}\beta_{i}\sigma\left(\left\langle\alpha_{i},x\right\rangle\right),$ (1)
where $\langle\alpha,x\rangle$ is the inner product $\alpha^{T}x$. We note that the total number of parameters in the GN is $nm+md$. As this count is proportional to the hidden layer size $m$, we use $m$ to characterize the model size.
General equivariant networks (GENs)
GENs represent a subset of GNs that constrain the output function $F$ to be equivalent with respect to a given group representation. A group representation for a group $G$ is formally defined as a homomorphism $\rho:G\to GL(m)$, where $GL(m)$ is the group of all invertible matrices on $\mathbb{R}^{m\times m}$. A function $F$ is considered equivariant if, for given input and output representations $\rho$ and $\phi$ respectively, it adheres to the condition $F\circ\rho_{g}=\phi_{g}\circ F$ for all group elements $g\in G$. A GN whose function $F$ satisfies this property is categorised as a GEN. Specifically, when $\phi=\text{id}$, the function $F$ is called an invariant function and the networks are called invariant networks. We note that GENs aim to represent all possible equivariant networks, and assume nothing but the equivariant output functions only.
Layer-wise equivariant networks (LENs)
LENs represent a subset of GNs that constrain each layer of the network to be equivalent with respect to a given group representation. Specifically, given the group representation $\psi$ in the hidden layer, LENs require that for all $g\in G$
$\phi_{g}\circ W^{(2)}=W^{(2)}\circ\psi_{g},~~~\psi_{g}\circ\sigma=\sigma\circ\psi_{g},~~~\text{and}~\psi_{g}\circ W^{(1)}=W^{(1)}\circ\rho_{g}.$
While each layer is constrained to be equivariant to maintain the symmetry in the data, the whole network is therefore equivariant. In fact, the output function is ensured to be equivariant under the group representations $\rho$ and $\phi$, because for every group element $g\in G$, we have
$F\circ\rho_{g}$ $=W^{(2)}\circ\sigma\circ W^{(1)}\circ\rho_{g}=W^{(2)}\circ\sigma\circ\psi_{g}\circ W^{(1)}$
$=W^{(2)}\circ\psi_{g}\circ\sigma\circ W^{(1)}=\phi_{g}\circ W^{(2)}\circ\sigma\circ W^{(1)}=\phi_{g}\circ F$
In practice, constructing LENs is a common approach to achieving GENs (he2021efficient). Different from GENs, LENs take more constraints. Firstly, one group representation $\psi_{g}$ making $\sigma\circ\psi_{g}=\psi_{g}\circ\sigma$ is needed for all $g\in G$, where we call $\psi$ admitted. Besides, both $W^{(1)}$ and $W^{(2)}$ should be equivalent under some representations, where we call a matrix in the intertwiner space. Given the group representations $\rho$ and $\phi$, to construct LENs is to find an admitted group representation $\psi$ and the corresponding intertwiner space.
Expressive power
In this paper, we focus on the expressive power of GNs, GENs, and LENs, which is defined as the minimum expected loss for approximating a target function $s$, and a smaller expected loss means a better expressive power. Since GENs serve as a subset of GNs and LENs form more constraints, trivially, we have for any target function $s$, any input distribution $P\in\Delta\mathbb{R}^{n}$, and any number of neurons $m$:
$\inf_{\theta\in\Theta_{m}\cap\text{GNs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}]\leq\inf_{\theta\in\Theta_{m}\cap\text{GENs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}]\leq\inf_{\theta\in\Theta_{m}\cap\text{LENs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}],$
where $\theta$ represents the parameter of a two-layer ReLU network ($\theta=(W^{(2)},W^{(1)})$), and $F_{\theta}$ is the forward function given parameter $\theta$, and $\Theta_{m}$ is the parameter set of $m$-hidden-layer networks. We reuse the notations GNs (also for GENs and LENs) to denote the respective parameter sets that makes $F_{\theta}$ a GN (also for GEN and LEN). To question whether equivariance limits the expressive power of two-layer network is to question whether there exists some equivariant function $s$ and $m$ such that
$\inf_{\theta\in\Theta_{m}\cap\text{GNs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}]$ $<\inf_{\theta\in\Theta_{m}\cap\text{GENs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}]$
$\text{and}~~~~\inf_{\theta\in\Theta_{m}\cap\text{GENs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}]$ $<\inf_{\theta\in\Theta_{m}\cap\text{LENs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}].$
For the first part, it questions whether equivariance is necessary for being the optimal approximator if the target function is equivariant. As for the second part, the question becomes whether layer-wise equivariance is the optimal way of incorporating equivariance. We also note that it is reasonable to assume the target function $s$ to be equivariant, because equivariant networks are designed for equivariant tasks.
4 Symmetric Boundary Hyperplanes and Symmetric Channel Vectors
To establish a clear foundation for our analysis of expressive power, we define two new terms, symmetry boundary hyperplanes and symmetry channel vectors.
4.1 GENs implies Symmetric Boundary Hyperplanes
Since the ReLU non-linearity is a piece-wise linear function, neural networks with only ReLU as activity function are also a piece-wise linear function. Specifically, we have the following formula for two-layer networks:
$F(x)=\sum_{i=1}^{m}\beta_{i}\sigma\left(\left\langle\alpha_{i},x\right\rangle\right)=\sum_{i:\left\langle\alpha_{i},x\right\rangle\geq 0}^{m}\beta_{i}\alpha_{i}^{\top}x=\left\langle\sum_{i:\left\langle\alpha_{i},x\right\rangle\geq 0}^{m}\alpha_{i}\beta_{i}^{\top},x\right\rangle$ $:=\left\langle\mathcal{F}(x),x\right\rangle,$ (2)
where $\mathcal{F}(x)=\sum_{i:\left\langle\alpha_{i},x\right\rangle\geq 0}^{m}\alpha_{i}\beta_{i}^{\top}$ is called the feature function of $F$ and is piece-wise constant.
We define the boundary hyperplane as the hyperplane whose feature function values are almost surely different on its two distinct sides. For simplicity, we assume that all hyperplanes contain 0 and have dimension n1n-1. The formal definition is shown below.
Figure 2: A visualization of the feature function $\mathcal{F}$ of $F(x,y)=\sigma(x)+\sigma(-y)+\sigma(-x+y)$, where the left subfigure is $\mathcal{F}_{1}$ and the right figure is $\mathcal{F}_{2}$. As shown, there are two boundary hyperplanes: $x=0$, $y=0$, $y=x$.
Definition 4.1 (boundary hyperplane).
Let $F$ be an output function of a two-layer ReLU network and $\mathcal{F}(x)$ be the corresponding feature function. We call a hyperplane $M$ a boundary hyperplane if and only if for all $x\in M$ and $v\not\in M$:
$\lim_{\delta\to 0^{+}}\mathcal{F}(x+\delta v)\not=\lim_{\delta\to 0^{+}}\mathcal{F}(x-\delta v).$
It is worth noting that we have no necessary to define the feature function $\mathcal{F}$ on the boundary hyperplanes. For any given function $F$, we can define it only on the set $S=\{x:F\text{ is linear in }B(x,r),\exists r>0\}$, where $B(x,r)=\{x+z:\|z\|<r\}$ is the open ball centering at $x$ with radius $r$. The value of $\mathcal{F}(x)$ is determined uniquely by the Riesz representation theorem. The complementary set $S^{c}$ is the union of all boundary hyperplanes, whose intersection with the hyperplanes $M\pm\delta v$.
For all $x\in S$ that satisfies $\langle x,\alpha_{i}\rangle\not=0$ for all $i\in[m]$, $\mathcal{F}(x)$ must be $\sum_{i:\langle x,\alpha_{i}\rangle>0}\beta_{i}\alpha_{i}^{T}$. Therefore, if the feature function changes when crossing through same hyperplane, the active set $\{i:\langle x,\alpha_{i}\rangle>0\}$ must change and the increment of the feature function is $\sum_{i:\langle\alpha_{i},x\rangle=0}\beta_{i}\alpha_{i}^{T}$, which is then summarized as Theorem 4.2, which can be viewed as an equivalent definition by the channel vectors.
Theorem 4.2.
Let $F=\sum_{i=1}^{m}\beta_{i}\sigma(\langle\alpha_{i},x\rangle)$ be an output function. Then, a hyperplane $M$ is a boundary hyperplane if and only for all $v\not\in M$:
$\sum_{i:\langle\alpha_{i},v\rangle>0,\langle\alpha_{i},M\rangle=0}\alpha_{i}\beta_{i}^{T}\not=\sum_{i:\langle\alpha_{i},v\rangle<0,\langle\alpha_{i},M\rangle=0}\alpha_{i}\beta_{i}^{T},$
where $\langle\alpha_{i},M\rangle=0$ are defined as $\langle\alpha_{i},x\rangle=0$ for all $x\in M$. Especially, if $M$ is a boundary hyperplane, there exists some channel vector $\alpha$ such that $\langle\alpha,M\rangle=0$.
Boundary hyperplane provides a powerful analytical tool for understanding the internal mechanisms of GENs. We prove that for a GEN, the set of its boundary hyperplanes are symmetric with respect to all group representations. Specifically, for any given boundary hyperplane $M$ and all fixed group representation $\rho$, applying the transformation $\rho$ to the boundary hyperplane (denoted as $\rho M=\{\rho x:x\in M\}$) also gets a boundary hyperplane, because:
$\lim_{\delta\to 0^{+}}\mathcal{F}(\rho x+\delta\rho v)=(\rho^{T})^{-1}\lim_{\delta\to 0^{+}}\mathcal{F}(x+\delta v)~\neq~(\rho^{T})^{-1}\lim_{\delta\to 0^{+}}\mathcal{F}(x-\delta v)=\lim_{\delta\to 0^{+}}\mathcal{F}(\rho x-\delta\rho v),$
where $x$ is a point on the hyperplane $M$ and $v$ is any vector not in $M$. The inequality holds because the group representation $\rho$ is always invertible. We note that $\rho M$ is not always another hyperplane. In fact, for the identity $e$, $\rho_{e}M=M$.
This property hold for all group elements $g\in G$, implying the set of all transformed hyperplanes, $\{\rho_{g}M:g\in G\}$, is a subset of the GEN’s total set of boundary hyperplanes. It also holds for all boundary hyperplanes $M$, which means the GEN’s total set of boundary hyperplanes is symmetry under $\rho$. We summarize this finding in the following theorem.
Theorem 4.3.
For a GEN on group $G$, its boundary hyperplane set $\mathcal{M}$ is closed under group representation $\rho_{g}$ for all $G$.
This theorem has direct implications for the expressive power of GENs. When approximating a target function, the network must capture its discontinuous, i.e., boundary hyperplanes. Due to the symmetry requirement, even a single learned hyperplane $M$ generates a whole family of hyperplanes, $\{\rho_{g}M:g\in G\}$. The size of this set provides a lower bound on the number of boundary hyperplanes the GEN must use to approximate the target function effectively. According to Theorem 4.2, a larger number of required hyperplanes, in turn, necessitates a greater number of neurons in the network.
4.2 LENs further implies Symmetric Channel Vectors
Compared with GENs, LENs enforce the symmetry constraints on every layer of the networks. While the hyperplane symmetry previously discussed is a property of all GENs, these further constraints imposed by the LEN structure lead to additional insights. Specifically, as defined before, LENs require one admitted group representation $\psi$ for ReLU nonlinearity and two matrices in the corresponding intertwining space.
We firstly consider the property of admitted group representation, i.e., $\psi\circ\sigma=\sigma\circ\psi$. This commutation property is not trivial and places a significant restriction on the form that the group representation $\psi$ can take. Specifically, for $\sigma=\operatorname{ReLU}$, the following lemma provides a complete characterization of all such admitted representations.
Lemma 4.4.
For $\sigma=\operatorname{ReLU}$, a group representation $\psi$ is admitted if and only if for all $g\in G$, $\psi_{g}$ is a generalized permutation matrix with exclusively positive non-zero entries. That is, each row and each column $\psi_{g}$ must contain exactly one non-zero element, and that element must be positive.
Proof.
We first prove that each row of $\psi_{g}$ has at most one non-zero entry for all $g\in G$. Assume for contradiction that exists $g\in G$ where a row $i$ of $\psi_{g}$ has at least two non-zero entries, $(\psi_{g})_{ij}$ and $(\psi_{g})_{ik}$. Without loss of generality, assume $(\psi_{g})_{ij}\leq(\psi_{g})_{ik}$. Input vector $x=e_{j}-e_{k}$ gets $(\psi_{g}\sigma(x))_{i}=(\psi_{g}e_{j})_{i}=(\psi_{g})_{ij}$ while $(\sigma(\psi_{g}x))_{i}=\sigma({(\psi_{g})_{ij}-(\psi_{g})_{ik}})=0$. Admitted $\psi$ implies $(\psi_{g})_{ij}=0$, which contradicts our initial assumption that the entries were non-zero. Therefore, each row of $\psi_{g}$ can contain at most one non-zero entry.
Each row has exactly one non-zero entry because $\psi_{g}$ is invertible, which also means each entry has at least one non-zero row, making $\psi_{g}$ a generalized permutation matrix. Moreover, when $(\psi_{g})_{ij}\neq 0$, we can set $x=e_{j}$. Then $\sigma(\psi_{g}x)=\psi_{g}\sigma(x)$ implies that $\sigma(\psi_{ij})=\psi_{ij}$. From the definition of ReLU, $(\psi_{g})_{ij}$ is positive for all $i$. ∎
Lemma 4.4 implies that the action of any $\psi_{g}$ on a vector is a composition of a permutation and a positive scaling of its coordinates. For each $g\in G$, we can identify a unique permutation $P_{g}$ that maps the coordinate indices. Specifically, $P_{g}(i)$ is the index of the only non-zero element in row $i$ of matrix $\psi_{g}$. A special case of an admitted group representation occurs when all non-zero entries of $\psi_{g}$ is 1. In this scenario, each $\psi_{g}$ is a standard permutation matrix. We refer to this $\psi$ as a permuting representation.
Following the property of admitted representation, we now examine how it interacts with the other constraints of invariant LENs. The final linear layer, represented by $W^{(2)}=[\beta_{1},\cdots,\beta_{m}]$, is required to satisfy the condition $W^{(2)}\circ\psi=W^{(2)}$ for LENs. This requirement reveals a deeper constraint on $\psi$ and the geometry of $W^{(2)}$. Specifically, we have the following lemma:
Lemma 4.5.
For admitted $\psi$ in an invariant LEN and any connected pair of indices $(i,j)$ (exists $g\in G$ making $P_{g}(i)=j$), we have $\|\beta_{j}\|\beta_{i}=\|\beta_{i}\|\beta_{j}$. Besides, all $h\in G$ with $P_{h}(i)=j$ shares the same positive $(\psi_{h})_{ij}$.
Proof.
Consider for a pair $(i,j)$ making $\psi_{ij}\neq 0$ for some $g\neq h$. Input vector $e_{j}$ gets $\beta_{j}=W^{(2)}e_{j}=W^{(2)}\psi e_{j}=\psi_{ij}W^{(2)}e_{i}=\psi_{ij}\beta_{i}$ for $\psi=\psi_{g}$ and $\psi=\psi_{h}$. Considering the direction of both sides, it shows that $\beta_{i}$ and $\beta_{j}$ has the same direction, meaning $\|\beta_{j}\|\beta_{i}=\|\beta_{i}\|\beta_{j}$. Considering the magnitude, the $\psi_{ij}$ does not change if $\psi$ switches from $\psi_{g}$ to $\psi_{h}$, meaning $(\psi_{g})_{ij}=(\psi_{h})_{ij}$. ∎
In light of this geometric alignment, we define that $\beta_{i}$ and $\beta_{j}$ are in the same orbit if they have the same direction, i.e., $\|\beta_{j}\|\beta_{i}=\|\beta_{i}\|\beta_{j}$. The lemma demonstrates that the invariance constraint forces all weight vectors within the same orbit to share the same direction in vector space. This imposes a strong, predictable structure on the weight space of the network’s final layer, partitioning it into sets of aligned vectors.
Aside from the constraint on $W^{(2)}$, LENs also impose a condition on $W^{(1)}$, which should belong to the intertwiner space defined by $\psi$ and $\rho$. Combined with the constraints on $\psi$ and $W^{(2)}$, we provide below a remarkable structural property for LENs.
Theorem 4.6.
For any $\rho$-invariant LEN $F_{\theta}$, there exists an equivalent $\rho$-invariant LEN, $F_{\tilde{\theta}}$ of the same size that computes the exact same function ($F_{\tilde{\theta}}(x)=F_{\theta}(x)$ for all $x$), where the channel vectors set of the rewritten network ($\{\tilde{\alpha}_{i}\}$) is closed under the action of the transposed group representation $\rho^{\top}$. That is, for any channel vector $\tilde{\alpha}_{i}$ and any group element $g\in G$, $\rho_{g}^{\top}\tilde{\alpha}_{i}$ is also a channel vector of $F_{\tilde{\theta}}$.
The proof is included in Appendix D. The theorem means that we can rewrite any LEN into another LEN of the same size, where the channel vectors of the rewritten LEN exhibit an explicit symmetry in $\rho^{\top}$. In the view of input-output mapping, the rewritten LEN is equivalent to the given LEN. For the rewritten LEN, the requirement of symmetry channel vectors means that one single channel vector $\tilde{\alpha}_{i}$ implicitly defines a subset of channel vectors $\{\rho_{g}^{\top}\tilde{\alpha}_{i}:g\in G\}$. The network size is then smaller and bounded to accommodate all these channel vectors.
We note that this constraint is stricter than the one identified for GENs. A GEN requires the symmetry for boundary hyperplanes while LEN requires that for channel vectors, where one boundary hyperplane has to be determined by at least one channel vector but different channel vectors could leads to one hyperplane. For instance, the channel vectors $\alpha$ and $\alpha$ are distinct vectors but define the exact same boundary hyperplane. A GEN has the flexibility to use scaled versions of the same vector normal, whereas a LEN, in its symmetric form, must treat each parallel vector as a distinct channel vector.
5 Drawback of Equivariance in Expressive Power
In this section, we show our results on how enforcing equivariance hurts the expressive power by constructing a target function on $\mathbb{R}^{2}\to\mathbb{R}$. When approximating this target function, we show that for neural networks with one neuron, i.e. in parameter set $\Theta_{1}$, GNs have strictly smaller expected loss than GENs. For 3-neuron-networks, GNs and GENs can fully approximate this target function, while LENs can not.
Example 5.1.
Consider group $G$ consists of 4 elements $e,g,h,gh$ where $g^{2}=h^{2}=e$ and $g\cdot h=h\cdot g=gh$. One group representation $\rho$ is as follows
$\rho_{e}=\begin{bmatrix}1&0\\ 0&1\\ \end{bmatrix},~~~~\rho_{g}=\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix},~~~~\rho_{h}=\begin{bmatrix}-1&0\\ 0&-1\\ \end{bmatrix},~~\text{and}~~\rho_{gh}=\begin{bmatrix}0&-1\\ -1&0\\ \end{bmatrix}$.
Given representation $\rho$, consider target function $s$:
$s:\mathbb{R}^{2}\to\mathbb{R},~~~~~(a,b)^{\top}\to\max(|a|,|b|,|a-b|),$
which is an invariant target function because
$s\circ\rho_{g}\begin{pmatrix}a\\ b\end{pmatrix}=s\begin{pmatrix}b\\ a\end{pmatrix}=\max(|b|,|a|,|b-a|)=s\begin{pmatrix}a\\ b\end{pmatrix}$,
$s\circ\rho_{h}\begin{pmatrix}a\\ b\end{pmatrix}=s\begin{pmatrix}-a\\ -b\end{pmatrix}=\max(|-a|,|-b|,|b-a|)=s\begin{pmatrix}a\\ b\end{pmatrix}$.
We note that $s$ has another form: $s((a,b)^{\top})=\sigma(a)+\sigma(-b)+\sigma(b-a)$.
Figure 1 provides a 3D plot and a 2D contour map to visualize the effects of the imposed symmetry. Furthermore, Figure 2 displays the feature function of the symmetry group element sss. Building on this example, we demonstrate how the constraints of GENs can limit expressive power and how this limitation is further pronounced in LENs. Specifically, we mainly focus on GENs vs GNs on the 1-neuron setting and LENs vs GENs on the 3-neuron setting.
For one-neuron networks
the output function of the neural network becomes $\beta\sigma(\langle\alpha,x\rangle)$. We consider GNs first, where no equivariance is required on the network. One trivial one-neuron network is the ReLU output of every dimension of input. Formally, when $\alpha=(1,0)^{\top}$ and $\beta=1$, we have $F((a,b)^{\top})=\sigma(a)$ being an approximation and $F$ is a GN with one neuron. Similarly when $\alpha=(0,1)^{\top}$ and $\beta=1$, $F((a,b)^{\top})=\sigma(b)$ is also a GN approximation. Thus, for all input distribution $P$ with $\operatorname{Pr}_{x\sim P}[x^{\top}=(0,0)]<1$, we have
$\inf_{\theta\in\Theta_{1}\cap\text{GNs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}]\leq\min\left[\mathbb{E}_{x\sim P}\|\sigma(a)-s(x)\|^{2},\mathbb{E}_{x\sim P}\|\sigma(b)-s(x)\|^{2}\right]<\mathbb{E}_{x\sim P}\|s(x)\|^{2}.$ (3)
For GENs, which has property $f(\rho x)=f(x)$ for $f(x)=\langle\mathcal{F}(x),x\rangle$ and $\rho$ being $\rho_{g}$ and $\rho_{h}$. Then we have $\langle\mathcal{F}(\rho x),\rho x\rangle=\langle\rho^{T}\mathcal{F}(\rho x),x\rangle=\langle\mathcal{F}(x),x\rangle$, which implies $\rho^{T}\circ\mathcal{F}\circ\rho=\mathcal{F}$. As a result, we have for all $x$:
$\beta\sigma(\langle\alpha,x\rangle)=\mathcal{F}(x)=(\rho_{h}^{T}\circ\mathcal{F}\circ\rho_{h})(x)=-\beta\sigma(\langle\alpha,-x\rangle),$
where at least one of $\sigma(\langle\alpha,x\rangle)$ and $\sigma(\langle\alpha,-x\rangle)$ is zero, indicating $\mathcal{F}=0$ for all $x$. It shows that the only one-neuron GEN is zero, which leads to strictly less expressive power then GNs. Specifically, we have for all input distribution $P$ with $\operatorname{Pr}_{x\sim P}[x^{\top}=(0,0)]<1$:
$\inf_{\theta\in\Theta_{1}\cap\text{GENs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}]=\mathbb{E}_{x\sim P}\|s(x)\|^{2}>\inf_{\theta\in\Theta_{1}\cap\text{GNs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}].$ (4)
For three-neurons networks
GNs can represent $s$ defined in Example 5.1 directly because the form $s((a,b)^{\top})=\sigma(a)+\sigma(-b)+\sigma(b-a)$ naturally fits the form of a three-neuron network. Due to the equivariance property of $s$, this specific network is also a GEN. Therefore, 3 neurons are sufficient for both GNs and GENs to fully approximate $s$.
However, for LENs, we show that 3 neurons are not enough; in fact, LENs require at least 6 neurons to fully approximate $s$. According to Theorem 4.6, for any LEN that approximates $s$, there exists a rewritten LEN with symmetric channel vectors of the same size that also perfectly approximates $s$. Due to the three boundary hyperplanes of $s$, this rewritten LEN must define three hyperplanes: $x=0$, $y=0$, and xy=0x-y=0, which implies the existence of at least three channel vectors: $c_{1}(1,0)$, $c_{2}(0,1)$, and $c_{3}(1,1)$ with $c_{1}c_{2}c_{3}\neq 0$. Since the channel vectors are symmetric, we have $\rho_{h}^{\top}c_{1}(1,0)=-c_{1}(1,0)$ is also a channel vector, so do $c_{2}(0,1)$ and $c_{3}(1,1)$. Consequently, the origin LEN must also contain at least 6 channel vectors, which applies to any LEN that fully approximates $s$.
Finally, we confirm that 6 neurons are sufficient, as the following LEN fully approximates $s$:
$F_{\theta}((a,b)^{\top})=\frac{1}{2}[\sigma(a)+\sigma(-a)+\sigma(b)+\sigma(-b)+\sigma(b-a)+\sigma(a-b)]$
6 Compensation to Enforcing Equivariance
In the previous sections, we demonstrated that for a fixed model size, the expressive power of LENs may strictly less than that of GENs, which are also strictly less expressive than GNs. A natural question arises: can this limitation in expressive power be overcome by increasing the model size? This section affirms that increasing model size is indeed a viable solution. We will examine this through both realizable and non-realizable cases. Furthermore, we also show that a key advantage of equivariant networks is their reduced hypothesis complexity, which leads to better generalizability.
6.1 Realizable case: Enlarging model size of LENs to compensate expressive power
We first consider the realizable case, where the target function can be perfectly represented by a GN with a finite number of neurons. By definition, this also implies that $s$ can be represented by a finite-neuron GEN. Our objective is to demonstrate the ability of LEN to approximate such a finite-neuron GEN.
Firstly, we show that with a larger model size, there always exists a LEN that can represent any representable mapping of finite-neuron GENs. To demonstrate this, consider a target GEN function defined as $s(x)=\sum_{i=1}^{m}\beta_{i}\sigma(\langle\alpha_{i},x\rangle)$. We can construct a LEN, denoted by $F(x)$, that perfectly replicates $s(x)$. The weights for this constructed LEN are defined as follows:
$W^{(2)}=\frac{1}{|G|}[\underbrace{\beta_{1},\dots,\beta_{1}}_{|G|},\dots,\underbrace{\beta_{m},\dots,\beta_{m}}_{|G|}]~~~~~\text{and}$
$W^{(1)}=[\rho_{g_{1}}^{T}\alpha_{1},\dots,\rho_{g_{|G|}}^{T}\alpha_{1},\dots,\rho_{g_{1}}^{T}\alpha_{m},\dots,\rho_{g_{|G|}}^{T}\alpha_{m}]^{T}.$
This construction ensures that the weights adhere to the necessary equivariance conditions. For the first layer, we have for all $g\in G$, $\rho_{g_{i}}\left\langle\alpha_{k},\rho_{g_{i}}(x)\right\rangle=\left\langle\rho_{g_{i}}^{\top}\rho_{g_{i}}^{\top}\alpha_{k},x\right\rangle=\left\langle\rho_{g_{i}g_{j}}^{\top}\alpha_{k},x\right\rangle$, which implies the existence of a permutation $\psi$ that maps the neuron index corresponding to $\rho_{g_{i}}^{\top}\alpha_{k}$ to the index for $\rho_{g_{i}g_{j}}^{\top}\alpha_{k}$. This $\psi$ is admitted because of Theorem 4.4. For $W^{(2)}$, we have the $\beta$ for all $\rho_{h}^{\top}\alpha_{i}$ is always $\beta$, indicating its invariance with respect to $\psi$.
This constructed LEN is a rewritten version of $s$ because of its invariance and equivariance. Specifically, we have
$F(x)$ $=\sum_{i=1}^{m}\frac{1}{|G|}\beta_{i}\sum_{g\in G}\sigma\left(\left\langle\rho_{g}^{\top}\alpha_{i},x\right\rangle\right)$
$=\frac{1}{|G|}\sum_{g\in G}\sum_{i=1}^{m}\beta_{i}\sigma\left(\left\langle\alpha_{i},\rho_{g}x\right\rangle\right)$
$=\frac{1}{|G|}\sum_{g\in G}s(\rho_{g}(x))=s(x),$
which confirms that our constructed LEN perfectly represents the target function $s$. This principle can also be extended to handle equivariant functions and multi-layer architectures, as detailed in Appendix E.
Building on this, we utilize the relationship between the number of boundary hyperplanes and that of channel vectors to study how many neurons are required to let LEN approach GEN. For GENs/LENs, the number of channel vectors are larger than the number of boundary hyperplanes by Theorem 4.2. Inversely for GENs, if there are $k$ boundary hyperplanes of some output function $k$, we can make some specific $\beta_{i}=1$ and $\alpha_{i}$ for $i\in[k]$ such that $f-\sum_{i=1}^{k}\beta_{i}\sigma(\langle\alpha_{i},x\rangle)$ has no boundary hyperplanes, which is then can be represented as a linear function $\langle\alpha_{0},x\rangle=\sigma(\langle\alpha_{0},x\rangle)-\sigma(\langle-\alpha_{0},x\rangle)$. Therefore, the number of boundary hyperplanes is larger than the number of channel vectors plus two for GENs. As for LENs, we utilize the following lemma to deal with all symmetric channel vectors as a whole part.
Lemma 6.1.
Let $\sum_{i=1}^{\ell}\beta_{i}\sum_{g\in G}\sigma(\langle\rho_{g}^{T}\alpha_{i},x\rangle)$ be an output function of LENs with $\alpha_{i}\not=0$ for all $i\in[\ell]$. Denote $M_{i}=\{\rho_{g}^{T}\alpha_{i}/\|\rho_{g}^{T}\alpha_{i}\|:g\in G\}$ as the normalized orbit induced by $\alpha_{i}$ and then we have $M_{i}=M_{j}$ or $M_{i}\cap M_{j}=\phi$ for all $i,j\in[\ell]$. Moreover, if $M_{i}=M_{j}$, we have
$\sum_{g\in G}\sigma(\langle\rho_{g}^{T}\alpha_{i},x\rangle)=k_{ij}\sum_{g\in G}\sigma(\langle\rho_{g}^{T}\alpha_{j},x\rangle),$
which enables us to reformulate the output function as $\sum_{i=1}^{\ell^{\prime}}\tilde{\beta_{i}}\sum_{g\in G}\sigma(\langle\rho_{g}^{T}\tilde{\alpha_{i}},x\rangle)$ such that $\tilde{M}_{i}\cap\tilde{M}_{j}=\phi$ for all $i,j\in[\ell^{\prime}]$. Furthermore, for each $\tilde{M}_{i}$, we can then merge all parallel channel vectors since
$\sum_{g\in G}\sigma(\langle\rho_{g}^{T}\tilde{\alpha}_{i},x\rangle)=|\text{Stab}(\tilde{\alpha}_{i})|\sum_{g\in G/\text{Stab}(\tilde{\alpha}_{i})}\sigma(\langle\rho_{g}^{T}\tilde{\alpha}_{i},x\rangle),$
where $\text{Stab}(\alpha)=\{g\in G:\rho_{g}^{T}\alpha=\alpha\}$ is the stabilizer subgroup with respected to $\alpha$. As a result, in the equivalent form $\sum_{i=1}^{\ell^{\prime}}\tilde{\beta}_{i}|\text{Stab}(\tilde{\alpha}_{i})|\sum_{g\in G/\text{Stab}(\tilde{\alpha}_{i})}\sigma(\langle\rho_{g}^{T}\tilde{\alpha}_{i},x\rangle)$, there are not two channel vectors of the same direction.
According to Lemma 6.1, we can construct LENs whose channel vectors are not parallel. Hence, the number of channel vectors is no more than double the number of boundary hyperplanes plus two. Therefore, the minimum required model size of LENs is approximately no more than double that of GENs, which is summarized in the following theorem with some technical modifications.
Theorem 6.2.
LENs require at most double model size to represent all invariant output functions of GENs.
This result indicates that enforcing layer-wise equivariance would be competitive among all methods of incorporating equivariance.
6.2 Non-realizable case: Enlarging model size can lead to even better expressive power
We now consider the non-realizable problem. With the same model size, equivariance may hurt the expressive power. We prove that when the networks have a larger model size, LENs may achieve a comparable and even better expressive power. We start with the following theorem, which shows a comparable expressive power when the model size is enlarged $|G|$ times.
Theorem 6.3.
Let $s$ be any invariant target function, and $\mathcal{Q}f$ the projection of $f$ to the equivariant space that is defined as $\mathcal{Q}f=\frac{1}{|G|}\sum_{g\in G}f\circ\rho_{g}$. If the data distribution is symmetric, then we have that
$\mathbb{E}[\|\mathcal{Q}f(x)-s(x)\|_{2}^{2}]\leq\mathbb{E}[\|{f}(x)-s(x)\|_{2}^{2}],$
for all approximator $f$. Specifically, if $f$ is an output function of GNs, we can find that $\mathcal{Q}f$ can be an output function of GENs/LENs with a $|G|$ times model size, implying that GENs/LENs can fix the expressive power gap and achieve a smaller expected loss by a larger model size.
When the GNs are of the output function $f(x)=\sum_{i=1}^{n}\beta_{i}\sigma(\langle\alpha_{i},x\rangle)$, the weight matrices of GENs/LENs can be formulated as
$W^{(2)}=\frac{1}{|G|}[\underbrace{\beta_{1},\dots,\beta_{1}}_{|G|},\dots,\underbrace{\beta_{m},\dots,\beta_{m}}_{|G|}]~\text{and}$
$W^{(1)}=[\rho_{g_{1}}^{T}\alpha_{1},\dots,\rho_{g_{|G|}}^{T}\alpha_{1},\dots,\rho_{g_{1}}^{T}\alpha_{m},\dots,\rho_{g_{|G|}}^{T}\alpha_{m}]^{T},$
which implies a comparable expressive power but a $|G|$ times model size. Meanwhile, the weight matrices can be generated in two steps: (1) copy the same number of original channel vectors, and (2) generate all symmetric channel vectors to make the network equivariant. Therefore, although LENs have a larger model size, they have the same number of learnable parameters. As a result, the dimension of the hypothesis space can be decreased by choosing less original channel vectors or constrain the original channel vectors in a subspace, while ensuring a comparable expressive power.
If we use the method in Lemma 6.1 to compress the networks, we can further get a better result. More precisely, we can let the weight matrices be
$W^{(2)}=[\underbrace{\beta_{1},\dots,\beta_{1}}_{C_{1}},\dots,\underbrace{\beta_{m},\dots,\beta_{m}}_{C_{m}}]~\text{and}$
$W^{(1)}=[\rho_{g_{11}}^{T}\alpha_{1},\dots,\rho_{g_{1C_{1}}}^{T}\alpha_{1},\dots,\rho_{g_{m1}}^{T}\alpha_{m},\dots,\rho_{g_{mC_{m}}}^{T}\alpha_{m}]^{T},$
where $\alpha_{i}$ is constrained to satisfy $|\{\rho_{g}^{T}\alpha_{i}:g\in G\}|=C_{i}$ and $g_{ij}$} is chosen to satisfy $\{\rho_{g_{i,j}}^{T}\alpha_{i}:j\in[C_{i}]\}=\{\rho_{g}^{T}\alpha_{i}:g\in G\}$. Similarly, this model still has comparable expressive power, but a lower-dimensional hypothesis space since we constrain all first-layer channel vectors in a (sub-)space.
Revise Example 5.1 where $C_{1}=4$ with arbitrary channel vector $(c_{1},c_{2})$ and $C_{2}=2$ with constrained channel vector $c_{3}(1,-1)$ that satisfies $\rho_{gh}(1,-1)=(1,-1)$. In this case, only two channel vectors are used originally to generate symmetric channel vectors, compared with $3$ for GNs. Moreover, the second channel vector $c_{3}(1,-1)$ are constrained in a one-dimension space, which further decreases the dimension of hypothesis space. As a result, the total numbers of parameters become
$\underbrace{2}_{\alpha_{1}}+\underbrace{1}_{\alpha_{2}}+\underbrace{2}_{\beta}\text{v.s.}\underbrace{2}_{\alpha_{1}}+\underbrace{2}_{\alpha_{2}}+\underbrace{2}_{\alpha_{3}}+\underbrace{3}_{\beta}.$
6.3 Constructed LEN with better generalizability
While increasing model size allows LENs to achieve comparable expressive power to GNs, a common concern is that this may harm generalization. In this section, we revisit Example 5.1 to demonstrate that despite requiring more neurons, LENs can still maintain a less complex hypothesis space than GNs. This lower complexity is significant as it suggests that LENs can achieve better generalization.
As discussed in Section 5, a LEN with 6 neurons and a GN with 3 neurons can both represent the target function. With the permutation representation $\psi$ defined as:
$\psi_{g}=\begin{bmatrix}0&1&&&&\\ 1&0&&&&\\ &&0&1&&\\ &&1&0&&\\ &&&&0&1\\ &&&&1&0\\ \end{bmatrix},\psi_{h}=\begin{bmatrix}&&1&0&&\\ &&0&1&&\\ 1&0&&&&\\ 0&1&&&&\\ &&&&0&1\\ &&&&1&0\\ \end{bmatrix}$.
The output function of the LEN can be formulated as:
$b_{1}[\sigma(c_{1}x+c_{2}y)+\sigma(c_{2}x+c_{1}y)]$
$+$ $b_{1}[\sigma(-c_{1}x-c_{2}y)+\sigma(-c_{2}x-c_{1}y)]$
$+$ $b_{2}[\sigma(c_{3}x-c_{3}y)+\sigma(-c_{3}x+c_{3}y)].$
From this formulation, we can determine that the dimension of the hypothesis space for the LEN (with 6 neurons) is 5. However, a GN with 3 neurons is defined as $\sum_{i=1}^{3}\beta_{i}\sigma(c_{1}x+c_{2}y)$, which means the 9-dimensional hypothesis space of the GN hypothesis space.
The key insight here is that although LENs may have a larger model size to maintain the expressive power, their inherent structural constraints could lead to a more lower-dimensional hypothesis space. Consequently, this reduced complexity implies that LENs are expected to exhibit superior generalization capabilities compared to standard GNs in this scenario.
7 Conclusions
In this work, we investigated the theoretical properties of equivariant neural networks, specifically focusing on the 2-layer ReLU architectures. Through the construction of an example, we demonstrated a key drawback of equivariance: it strictly reduces expressive power. Nevertheless, we proved that this reduction is recoverable by increasing the model size. We also show that although the larger model size is required to maintain expressivity, it could retain a lower complexity for the resulting hypothesis space, thereby still offering a superior generalizability for equivariant networks.
Appendix A Proof of Theorem 4.2
We verify that the boundary hyperplane is well-defined. Firstly, we should check the definition of feature function $\mathcal{F}$. Given the function $F=\sum_{i=1}^{m}\beta_{i}\sigma(\langle\alpha_{i},x\rangle)$, we could consider the set $S=\{x\in\mathbb{R}^{n}:\langle\alpha_{i},x\rangle\not=0,\forall i\in[m]\}$. Of course its complementary set $S^{c}$ is the union of some hyperplanes $\bigcup_{i=1}^{m}M_{i}$, where $M_{i}=\{x:\langle\alpha_{i},x\rangle=0\}$. Then for any $x\in S$, one of $\langle\alpha_{i},x\rangle>0$ and $\langle\alpha_{i},x\rangle<0$ holds. For small enough $r$, $\langle\alpha_{i},x\rangle>0$ or $\langle\alpha_{i},x\rangle<0$ holds for all $y\in B(x,r)$. As a result, we have $B(x,r)\subset S$. While every $\langle\alpha_{i},y\rangle$ does not change its sign for $y\in B(x,r)$, the output function $F$ is linear in $B(x,r)$. Especially by Riesz representation theorem, there exists a unique vector $v$ such that $F(y)=\langle v,y\rangle$ on $B(x,r)$. Combining with that $F(x)=\sum_{i:\langle\alpha_{i},x\rangle>0}\beta_{i}\langle\alpha_{i},x\rangle$, we have $\mathcal{F}(x)=\sum_{i:\langle\alpha_{i},x\rangle>0}\alpha_{i}\beta_{i}^{T}$ for all $x\in S$. Moreover, for every hyperplane $M\not=M_{i}$ for all $i\in[m]$ and any continuous distribution $m$ on it, we have $x\in S$ almost surely.
For every $x\in M$ of any given hyperplane $M$ and $v\not\in M$, we have $\mathcal{F}(x+\delta v)$ is a constant when $\delta>0$ is small enough, since $x+\delta v$ would not pass through any hyperplanes $M_{i}$ as $\delta$ goes to 0. As a result, $\mathcal{F}(x+\delta v)$ converges to that constant uniformly, bounded by $\sum_{i=1}^{m}\|\alpha_{i}\beta_{i}^{T}\|$. Therefore, dominated convergence theorem ensures that $\lim\mathbb{E}[\mathcal{F}(x+\delta v)]=\mathbb{E}[\lim\mathcal{F}(x+\delta v)]$. Moreover, if $x+\delta v$ is not in the the hyperplanes $M_{i}$ (it almost surely holds), $x\in S$ implies $\lim\mathcal{F}(y+\delta v)$ is a constant for all $y\in B(x,r)\cap M$ for a small radius $r$. Since $\lim\mathcal{F}(x+\delta v)$ is piece-wise constant, the expectation is well-defined.
For given $M$, we can compare $\lim\mathcal{F}(x+\delta v)$ and $\lim\mathcal{F}(x-\delta v)$ immediately. If $x\not\in M_{i}$ for all $i\in[m]$, we have $x\in S$ and then $x+\delta v$ and $x-\delta v$ are both in $B(x,r)\subset S$ for small $\delta$, resulting in that $\mathcal{F}(x+\delta v)=\mathcal{F}(x-\delta v)$ for small $\delta>0$. As a result, if $M\not=M_{i}$ for all $i\in[m]$, $x$ is in $S$ and therefore $\mathcal{F}(x+\delta v)=\mathcal{F}(x-\delta v)$ almost surely, so that $\lim\mathbb{E}[\mathcal{F}(x+\delta v)]=\lim\mathbb{E}[\mathcal{F}(x-\delta v)]$. Otherwise, we can let $\tilde{F}=F-\sum_{i:\langle\alpha_{i},x\rangle=0,\forall x\in M}\beta_{i}\sigma(\langle\alpha_{i},x\rangle)$ and thus
$\lim\mathbb{E}[\tilde{\mathcal{F}}(x+\delta v)]=\lim\mathbb{E}[\tilde{\mathcal{F}}(x-\delta v)].$
Meanwhile, let G{G} denote $F-\tilde{F}=\sum_{i:\langle\alpha_{i},x\rangle=0,\forall x\in M}\beta_{i}\sigma(\langle\alpha_{i},x\rangle)$, we have
$\mathcal{G}(x+\delta v)-\mathcal{G}(x-\delta v)=\sum_{i:\langle\alpha_{i},v\rangle>0}\alpha_{i}\beta_{i}^{T}-\sum_{i:\langle\alpha_{i},v\rangle<0}\alpha_{i}\beta_{i}^{T},\forall x\in M.$
Therefore, we have
$\mathbb{E}[\mathcal{G}(x+\delta v)-\mathcal{G}(x-\delta v)]=\sum_{i:\langle\alpha_{i},v\rangle>0}\alpha_{i}\beta_{i}^{T}-\sum_{i:\langle\alpha_{i},v\rangle<0}\alpha_{i}\beta_{i}^{T}.$
As a result, the hyperplane $M$ is boundary hyperplane if and only if $\sum_{i:\langle\alpha_{i},v\rangle>0}\alpha_{i}\beta_{i}^{T}\not=\sum_{i:\langle\alpha_{i},v\rangle<0}\alpha_{i}\beta_{i}^{T}$.
Moreover, if no $\alpha_{i}$ satisfies that $\langle\alpha_{i},M\rangle=0$, the hyperplane $M$ is not a boundar hyperplane since $\sum_{i:\langle\alpha_{i},v\rangle>0}\alpha_{i}\beta_{i}^{T}=0=\sum_{i:\langle\alpha_{i},v\rangle<0}\alpha_{i}\beta_{i}^{T}$. Inversely, if $M$ is a boundary hyperplane, there exists some channel vector $\alpha$ such that $\langle x,M\rangle=0$.
The proof is completed. In addition, we would like to use the feature function to characterize the boundary hyperplane for future extension to multi-layer conditions, where the equivaENt condition would be more complex. However, the feature functions can be similarly defined and therefore the boundary can be shaped.
Appendix B Proof of Lemma 4.4
For simplicity, we denote $e_{i}\in\mathbb{R}^{n}$ as the vector $(0,\dots,1,\dots,0)^{T}$, where the $i$-th entry is $1$ and other entries are all 0. Since $\psi_{g}\circ\sigma=\sigma\circ\psi_{g}$, we have $\psi\sigma(x)=\sigma(\psi x)$ for all $x\in\mathbb{R}^{n}$.
If there are two nonzero entries $\psi_{ij}<\psi_{ik}$ in the same row of $\psi_{g}$, we can set $x=e_{j}-e_{k}$ and then we have $(\psi_{g}\sigma(x))_{i}=(\psi_{g}e_{j})_{i}=\psi_{ij}$ while $(\sigma(\psi_{g}x))_{i}=\sigma({\psi_{ij}-\psi_{ik}})=0$. That implies $\psi_{ij}=0$, which contradicts. Thus, each row has at most one nonzero entry. Hence, there are at most $n$ nonzero entries of $\psi_{g}$, which are in different rows.
Since $\psi_{g}$ is invertible, each row and each column of $\psi_{g}$ have at least one nonzero entry. It implies that there are at least $n$ nonzero entries of $\psi_{g}$. Thus, there are exactly $n$ nonzero entries of $\psi_{g}$ and $\psi_{g}$ is a generalized permutation matrix $\text{diag}(\lambda_{g}^{1},\dots,\lambda_{g}^{m})P_{g}$.
Moreover, when $\psi_{ij}\not=0$, we can set $x=e_{j}$. Then $\sigma(\psi x)=\psi\sigma(x)$ implies that $\sigma(\psi_{ij})=\psi_{ij}$. From the definition of ReLU, we have that $\psi_{ij}>0$. The proof is completed.
Appendix C Proof of Theorem 4.3
Let $F$ be an invariant output function that can be represented as $\sum_{i=1}^{m}\beta_{i}\sigma(\langle\alpha_{i},x\rangle)$. Since $F$ is invariant, it can be also reformulated as
$F=\mathcal{Q}F=\sum_{i=1}^{m}\frac{\beta_{i}}{|G|}\sum_{g\in G}\sigma(\langle\alpha_{i},\rho_{g}x\rangle).$
Then the feature function can be
$\mathcal{F}(x)=\frac{1}{|G|}\sum_{i,g:\langle\rho_{g}^{T}\alpha_{i},x\rangle>0}\rho_{g}^{T}\alpha_{i}\beta_{i}^{T}.$
Then, for every transformed point $\rho_{g}x$, we have
$\mathcal{F}(\rho_{g}x)=$ $\frac{1}{|G|}\sum_{i,h:\langle\rho_{h}^{T}\alpha_{i},\rho_{g}x\rangle>0}\rho_{h}^{T}\alpha_{i}\beta_{i}^{T}=\frac{1}{|G|}\sum_{i,h:\langle\rho_{hg}^{T}\alpha_{i},x\rangle>0}(\rho_{g}^{T})^{-1}\rho_{hg}^{T}\alpha_{i}\beta_{i}^{T}$
$=$ $(\rho_{g}^{T})^{-1}\frac{1}{|G|}\sum_{i,h:\langle\rho_{h}^{T}\alpha_{i},x\rangle>0}\rho_{h}^{T}\alpha_{i}\beta_{i}^{T}=(\rho_{g}^{T})^{-1}\mathcal{F}(x).$
As a result, $\mathcal{F}$ is somehow “equivariant". Moreover, let us consider the definition of boundary hyperplanes, we have
$\lim_{\delta\to 0^{+}}\mathbb{E}[\mathcal{F}(\rho_{g}x+\rho_{g}v)]=\lim_{\delta\to 0^{+}}\mathbb{E}[\mathcal{F}(\rho_{g}(x+v))]=(\rho_{g}^{T})^{-1}\lim_{\delta\to 0^{+}}\mathbb{E}[\mathcal{F}(x+v)],$
and
$\lim_{\delta\to 0^{+}}\mathbb{E}[\mathcal{F}(\rho_{g}x-\rho_{g}v)]=\lim_{\delta\to 0^{+}}\mathbb{E}[\mathcal{F}(\rho_{g}(x-v))]=(\rho_{g}^{T})^{-1}\lim_{\delta\to 0^{+}}\mathbb{E}[\mathcal{F}(x-v)],$
where $v\not\in M$ and $\rho_{g}v\not\in\rho_{g}M$. Therefore, $M$ is a boundary hyperplane if and only if $\rho_{g}M$ is a boundary hyperplane.
Besides, if we consider the equivaENt condition, $M$ is a boundary hyperplane if and only if
$\sum_{\begin{subarray}{c}i,h:\langle\rho_{h}^{T}\alpha_{i},v\rangle>0\\ \langle\rho_{h}^{T}\alpha_{i},M\rangle=0\end{subarray}}\rho_{h}^{T}\alpha_{i}\beta_{i}^{T}\not=\sum_{\begin{subarray}{c}i,h:\langle\rho_{h}^{T}\alpha_{i},v\rangle<0\\ \langle\rho_{h}^{T}\alpha_{i},M\rangle=0\end{subarray}}\rho_{h}^{T}\alpha_{i}\beta_{i}^{T}.$
By contrast, that for $\rho_{g}M$ is
$\sum_{\begin{subarray}{c}i,h:\langle\rho_{h}^{T}\alpha_{i},\rho_{g}v\rangle>0\\ \langle\rho_{h}^{T}\alpha_{i},\rho_{g}M\rangle=0\end{subarray}}\rho_{h}^{T}\alpha_{i}\beta_{i}^{T}\not=\sum_{\begin{subarray}{c}i,h:\langle\rho_{h}^{T}\alpha_{i},\rho_{g}v\rangle<0\\ \langle\rho_{h}^{T}\alpha_{i},\rho_{g}M\rangle=0\end{subarray}}\rho_{h}^{T}\alpha_{i}\beta_{i}^{T}.$
The LHS is equal to
$\sum_{\begin{subarray}{c}i,h:\langle\rho_{h}^{T}\alpha_{i},\rho_{g}v\rangle>0\\ \langle\rho_{h}^{T}\alpha_{i},\rho_{g}M\rangle=0\end{subarray}}\rho_{h}^{T}\alpha_{i}\beta_{i}^{T}=\sum_{\begin{subarray}{c}i,h:\langle\rho_{hg}^{T}\alpha_{i},v\rangle>0\\ \langle\rho_{hg}^{T}\alpha_{i},M\rangle=0\end{subarray}}(\rho_{g}^{T})^{-1}\rho_{hg}^{T}\alpha_{i}\beta_{i}^{T}=(\rho_{g}^{T})^{-1}\sum_{\begin{subarray}{c}i,h:\langle\rho_{h}^{T}\alpha_{i},v\rangle>0\\ \langle\rho_{h}^{T}\alpha_{i},M\rangle=0\end{subarray}}\rho_{h}^{T}\alpha_{i}\beta_{i}^{T},$
while the RHS is equal to
$\sum_{\begin{subarray}{c}i,h:\langle\rho_{h}^{T}\alpha_{i},\rho_{g}v\rangle<0\\ \langle\rho_{h}^{T}\alpha_{i},\rho_{g}M\rangle=0\end{subarray}}\rho_{h}^{T}\alpha_{i}\beta_{i}^{T}=\sum_{\begin{subarray}{c}i,h:\langle\rho_{hg}^{T}\alpha_{i},v\rangle<0\\ \langle\rho_{hg}^{T}\alpha_{i},M\rangle=0\end{subarray}}(\rho_{g}^{T})^{-1}\rho_{hg}^{T}\alpha_{i}\beta_{i}^{T}=(\rho_{g}^{T})^{-1}\sum_{\begin{subarray}{c}i,h:\langle\rho_{h}^{T}\alpha_{i},v\rangle<0\\ \langle\rho_{h}^{T}\alpha_{i},M\rangle=0\end{subarray}}\rho_{h}^{T}\alpha_{i}\beta_{i}^{T}.$
As a result, $M$ and $\rho_{g}M$ have the equivaENt conditions to be boundary hyperplanes, causing that $M$ is a boundary hyperplane if and only if $\rho_{g}M$ is a boundary hyperplane. The proof is completed.
Appendix D Proof of Theorem 4.6
Due to $W^{(1)}\circ\rho_{g}=\psi_{g}\circ W^{(1)}$, and thus we have
$W^{(1)}$ $=\frac{1}{|G|}\sum_{g\in G}\psi_{g}^{-1}\circ W^{(1)}\circ\rho_{g}.$
Then $F$ can be rewritten as:
$F(x)=$ $\begin{bmatrix}\beta_{1}\ldots\beta_{m}\end{bmatrix}\sigma\left(\begin{bmatrix}\alpha_{1}&\cdots&\alpha_{m}\end{bmatrix}^{\top}x\right)$
$=$ $\begin{bmatrix}\beta_{1}\ldots\beta_{m}\end{bmatrix}\sigma\left(\frac{1}{|G|}\sum_{g\in G}\left(\psi_{g}^{-1}\begin{bmatrix}\alpha_{1}&\cdots&\alpha_{m}\end{bmatrix}\right)^{\top}\rho_{g}x\right)$
$=$ $\begin{bmatrix}\beta_{1}\ldots\beta_{m}\end{bmatrix}\sigma\left(\frac{1}{|G|}\sum_{g\in G}\begin{bmatrix}\frac{\alpha_{P_{g}(1)}}{(\psi_{g})_{1,P_{g}(1)}}&\cdots&\frac{\alpha_{P_{g}(m)}}{(\psi_{g})_{m,P_{g}(m)}}\end{bmatrix}^{\top}\rho_{g}x\right)$
$=$ $\frac{1}{|G|}\sum_{i=1}^{m}\beta_{i}\sigma\left(\sum_{g\in G}\left(\frac{\alpha_{P_{g}(i)}}{(\psi_{g})_{i,P_{g}(i)}}\right)^{\top}\rho_{g}x\right)$
$=$ $\frac{1}{|G|}\sum_{i=1}^{m}\frac{\beta_{i}}{\|\beta_{i}\|}\sigma\left(\sum_{g\in G}\left(\frac{\|\beta_{i}\|}{(\psi_{g})_{i,P_{g}(i)}}\alpha_{P_{g}(i)}\right)^{\top}\rho_{g}x\right)$
$=$ $\sum_{i=1}^{m}\frac{\beta_{i}}{\|\beta_{i}\|}\sigma\left(\left(\sum_{g\in G}\frac{1}{|G|}\rho_{g}^{\top}\|\beta_{P_{g}(i)}\|\alpha_{P_{g}(i)}\right)^{\top}x\right).$
Denote $\tilde{\beta_{i}}=\frac{\beta_{i}}{\|\beta_{i}\|}$ and $\tilde{\alpha_{i}}=\sum_{g\in G}\frac{1}{|G|}\rho_{g}^{\top}\|\beta_{P_{g}(i)}\|\alpha_{P_{g}(i)}$, we get $F(x)=\sum_{i=1}^{m}\tilde{\beta_{i}}\sigma\langle\tilde{\alpha_{i}},x\rangle$.
For new parameters, we have
$\rho_{h}^{\top}\tilde{\alpha_{i}}$ $=\rho_{h}^{\top}\sum_{g\in G}\frac{1}{|G|}\rho_{g}^{\top}\beta_{P_{g}(i)}\alpha_{P_{g}(i)}$
$=\sum_{g\in G}\frac{1}{|G|}\rho_{gh}^{\top}\beta_{P_{g}(i)}\alpha_{P_{g}(i)}$
$=\sum_{g\in G}\frac{1}{|G|}\rho_{gh}^{\top}\beta_{P_{gh}(P_{h^{-1}}(i))}\alpha_{P_{g}(P_{h^{-1}}(i))}$
$=\tilde{\alpha}_{P_{h^{-1}}(i)},$
which is a satisfied rewritten LEN.
Appendix E Proof of ENs’ Universal Represetation Ability
We construct the following ENs to represent all equivariant output functions of GNs. Let the output function of GNs be $W^{(L)}\sigma(W^{(L-1)}\sigma(\dots\sigma(W^{(1)}x)\dots))$. Let the weight matrices of ENs be
$\tilde{W}^{(L)}=\frac{1}{|G|}(\underbrace{\phi_{g_{1}}^{-1}W^{(L)},\dots,\phi_{g_{|G|}}^{-1}W^{(L)}}_{|G|})$
$\tilde{W}^{(\ell)}=\text{diag}(\underbrace{W^{(\ell)},\dots,W^{(\ell)}}_{|G|}),\forall\ell=2,3,\dots,L-1$
$\tilde{W}^{(1)}=(\underbrace{(W^{(1)}\rho_{g_{1}})^{T},\dots,(W^{(1)}\rho_{g_{|G|}})^{T}}_{|G|})^{T},$
where the group representation $\psi$ in the intermediate layers is the corresponding permutation representation acting on each axis $e_{g_{i}}$ as $\psi_{g}e_{g_{i}}=e_{g_{i}g^{-1}}$. As a result, we have
$\tilde{W}^{(1)}\rho_{g}=\begin{bmatrix}W^{(1)}\rho_{g_{1}}\\ W^{(1)}\rho_{g_{2}}\\ \vdots\\ W^{(1)}\rho_{g_{|G|}}\end{bmatrix}\rho_{g}=\begin{bmatrix}W^{(1)}\rho_{g_{1}g}\\ W^{(1)}\rho_{g_{2}g}\\ \vdots\\ W^{(1)}\rho_{g_{|G|}g}\end{bmatrix}=\psi_{g}\begin{bmatrix}W^{(1)}\rho_{g_{1}}\\ W^{(1)}\rho_{g_{2}}\\ \vdots\\ W^{(1)}\rho_{g_{|G|}}\end{bmatrix}=\psi_{g}\tilde{W}^{(1)},$
$\tilde{W}^{(\ell)}\psi_{g}=\text{diag}(\underbrace{W^{(\ell)},\dots,W^{(\ell)}}_{|G|})\psi_{g}=\text{diag}(\underbrace{W^{(\ell)},\dots,W^{(\ell)}}_{|G|})=\psi_{g}\text{diag}(\underbrace{W^{(\ell)},\dots,W^{(\ell)}}_{|G|})=\psi_{g}\tilde{W}^{(\ell)},\forall\ell=2,3\dots,L-1$
$\tilde{W}^{(L)}\psi_{g}=\frac{1}{|G|}(\underbrace{\phi_{g_{1}}^{-1}W^{(L)},\dots,\phi_{g_{|G|}}^{-1}W^{(L)}}_{|G|})\psi_{g}=\frac{1}{|G|}(\underbrace{\phi_{g_{1}g^{-1}}^{-1}W^{(L)},\dots,\phi_{g_{|G|g^{-1}}}^{-1}W^{(L)}}_{|G|})=\frac{1}{|G|}\phi_{g}(\underbrace{\phi_{g_{1}}^{-1}W^{(L)},\dots,\phi_{g_{|G|}}^{-1}W^{(L)}}_{|G|})$
As a result, the constructed network has the same output function $\mathcal{Q}F$ and equivariant layers, renders it an EN. The proof is completed.
Appendix F Proof of Lemma 6.1
Denote $M=\{\rho_{g}^{T}\alpha/\|\rho_{g}^{T}\alpha\|:g\in G\}$ and $N=\{\rho_{g}^{T}\beta/\|\rho_{g}^{T}\beta\|:g\in G\}$, where $\alpha$ and $\beta$ are two nonzero vectors in $\mathbb{R}^{n}$. Then we prove that $M\cap N=\emptyset$ or $M=N$.
If $M\cap N\not$$\emptyset$, there exist group elements $g_{0},h_{0}\in G$ such that $\rho_{g_{0}}^{T}\alpha/\|\rho_{g_{0}}^{T}\alpha\|=\rho_{h_{0}}^{T}\beta/\|\rho_{h_{0}}^{T}\beta\|$. Then, for any element $\rho_{g}^{T}\alpha/\|\rho_{g}^{T}\alpha\|\in M$, there exists $\rho^{T}_{h_{0}g_{0}^{-1}g}\beta/\|\rho^{T}_{h_{0}g_{0}^{-1}g}\beta\|\in N$ such that
$\rho^{T}_{h_{0}g_{0}^{-1}g}\beta=\rho_{g}^{T}\rho_{g_{0}^{-1}}^{T}\rho^{T}_{h_{0}}\beta=\frac{\|\rho_{h_{0}}^{T}\beta\|}{\|\rho_{g_{0}}^{T}\alpha\|}\cdot\rho_{g}^{T}\alpha~~\text{and}~~\frac{\rho^{T}_{h_{0}g_{0}^{-1}g}\beta}{\|\rho^{T}_{h_{0}g_{0}^{-1}g}\beta\|}=\frac{\rho_{g}^{T}\alpha}{\|\rho_{g}^{T}\alpha\|}.$
It implies that $M\subset N$. Inversely, we can prove that $N\subset M$ in the same way. Hence, combining $M\subset N$ and $N\subset M$, we have $M=N$. Moreover, when denoting $\widetilde{M}=\{\rho_{g}^{T}\alpha:g\in G\}$, we can also prove that $|\widetilde{M}|=|M|$. If there exist $g,h\in G$ such that $\rho_{g}^{T}\alpha/\|\rho_{g}^{T}\alpha\|=\rho_{h}^{T}\alpha/\|\rho_{h}^{T}\alpha\|$, we have $\rho_{g}^{T}\alpha=c\rho_{h}^{T}\alpha$ with some positive number $c$ since they have the same direction. Then, we have $\rho_{gh^{-1}}^{T}\alpha=c\alpha$ and
$(\rho_{gh^{-1}}^{T})^{|G|}\alpha=(\rho_{gh^{-1}}^{T})^{|G|-1}c\alpha=(\rho_{gh^{-1}}^{T})^{|G|-2}c^{2}\alpha=\dots=c^{|G|}\alpha.$
Meanwhile, since $(gh^{-1})^{|G|}=e\in G$, we have $(\rho_{gh^{-1}}^{T})^{|G|}\alpha=\rho_{e}^{T}\alpha=\alpha$. It implies that $c^{|G|}=1$ and thus $c=1$. Finally, we have that $\rho_{g}^{T}\alpha=\rho_{h}^{T}\alpha\Leftrightarrow\frac{\rho^{T}_{g}\alpha}{\|\rho^{T}_{g}\alpha\|}=\frac{\rho_{h}^{T}\alpha}{\|\rho_{h}^{T}\alpha\|}$, which indicates that $|\widetilde{M}|=|M|$. Therefore, if $M=N$, we have $\widetilde{M}=k\widetilde{N}$ for some $k>0$.
In particular, if there exist two parallel channel vectors $\rho_{s}^{T}\alpha_{i}$ and $\rho_{t}^{T}\alpha_{j}$, we can assume $\rho_{s}^{T}\alpha_{i}=k\rho_{t}^{T}\alpha_{j}$ for some $k>0$ and then we have
$\mathcal{O}_{i}=\{\rho_{g}^{T}(\rho_{s}^{T}\alpha_{i}):g\in G\}=\{\rho_{g}^{T}(k\rho_{t}^{T}\alpha_{j}):g\in G\}=k\{\rho_{g}^{T}(\rho_{t}^{T}\alpha_{j}):g\in G\}=k\mathcal{O}_{j}.$
Moreover, ENs can merge these channels as
$\sum_{g\in G}\beta_{i}\sigma(\langle\rho_{g}^{T}\alpha_{i},x\rangle)+\beta_{j}\sigma(\langle\rho_{g}^{T}\alpha_{j},x\rangle)=\sum_{g\in G}(k\beta_{i}+\beta_{j})\sigma(\langle\rho_{g}^{T}\alpha_{j},x\rangle)$
As a result, we can find an EN of the same output function such that $\rho_{g}^{T}\alpha_{i}/\|\rho_{g}^{T}\alpha_{i}\|\not=\rho_{h}^{T}\alpha_{j}/\|\rho_{h}^{T}\alpha_{j}\|$ for all $i\not=j$.
Furthermore, we can define the stabilizer subgroup $Stab(\alpha)=\{g:\rho_{g}^{T}\alpha_{i}\}$ and therefore $G=Stab(\alpha)\oplus\tilde{M}$ with
$\sum_{g\in G}f(\rho_{g}^{T}\alpha)=\sum_{h\in G/\text{Stab}(\alpha)}\sum_{gh^{-1}\in\text{Stab}(\alpha)}f(\rho_{g}^{T}\alpha)=|\text{Stab}(\alpha)|\cdot\sum_{h\in G/\text{Stab}(\alpha)}f(\rho_{h}^{T}\alpha).$
The proof is completed.
Appendix G Proof of Theorem 6.2
We consider the following conditions with given invariant output function of GNs $F(x)=\sum_{i=1}^{m}\beta_{i}\sigma(\langle\alpha_{i},x\rangle$. Denote the number of boundary hyperplane of $F(x)$ by $N$.
If $N\leq m-1$, we can construct the EN of Lemma 6.1, where we let the output function be formulated as $\mathcal{Q}F(x)=\frac{1}{|G|}\sum_{i=1}^{m}\beta_{i}\sum_{g\in G}\sigma(\langle\rho_{g}^{T}\alpha_{i},x\rangle)$. Lemma 6.1 implies that the number of channel vectors $m^{\prime}$ after merging is at most double the number of boundary hyperplanes plus $2$, and therefore,
$m^{\prime}\leq 2N+2\leq 2(m-1)+2=2m.$
If $N=m$, there are $m$ boundary hyperplanes $M_{1},\dots,M_{m}$ and we can assume $\langle\alpha_{i},M_{i}\rangle=0$ for all $i\in[m]$ without loss of generality. Theorem 4.3 indicates that all the boundary hyperplanes are symmetric, enable us to classifier them as $S_{1}=\{M_{k_{0}+1},\dots,M_{k_{1}}\},\dots,S_{l+1}=\{M_{k_{l}+1},\dots,M_{k_{l+1}}\}$ with $k_{0}=0$ and $k_{l+1}=m$, where each set collects all symmetric boundary hyperplanes. Then consider the orbits $\mathcal{O}_{i}=\{\rho_{g}^{T}\alpha_{i}:g\in G\}$, we have $\mathcal{O}_{i}=k_{ij}\mathcal{O}_{j}$ for all $i,j$ in the same classification set $S$. Lemma 6.1 implies that the sum $\frac{1}{|G|}\sum_{i=1}^{m}\beta_{i}\sum_{g\in G}\sigma(\langle\rho_{g}^{T}\alpha_{i},x\rangle)$ can be formulated by the sum of at most $2m$ items, resulting in $m^{\prime}\leq 2m$ also.
The proof is completed.